OpenStudy (anonymous):

Solve this system algebraically. 7x - 2y = 4 5y + 3x = 10 a){(-41/40, -41/58)} b){(41/40, 41/58)} c){(-40/41, -58/41)} d){(40/41, 58/41)}

OpenStudy (anonymous):

-2y=4-7x y=(7x-4)/2 sub into 2nd equation 5((7x-4)/2)+3x=10 35x/2-20/2+3x=10 41x/2=20 41x=40 x=40/41 subs x=40/41 into eqtn 2 5y+3(40/41)=10 5y+120/41=10 205y+120=410 205y=290 y=290/205 which can be cancelled down to; y=58/41 So d is the right answer .. hope this is what u were hoping for .. sorry it's so long winded :)